F(x)=3x^2+4
Find f(3) with this part I got the answer 31
Find 3f(-x) this one is tricky for me
Plz help

Answer by jim_thompson5910(35256)


You have the correct value for f(3) since
f(x) = 3x^2+4
f(3) = 3(3)^2+4
f(3) = 3(9)+4
f(3) = 27+4
f(3) = 31
We replaced x with 3 and evaluated

--------------------------------------------------

A similar idea will happen with the next part. Let's first find f(-x)
So,
f(x) = 3x^2+4
f(-x) = 3(-x)^2+4 ... replace every x with -x
f(-x) = 3x^2+4 ... note how (-x)^2 = x^2
This is because the -x really means -1x, and we're squaring -1 to get +1 or just 1.

Now we multiply both sides by 3 to get...
f(-x) = 3x^2+4
3*f(-x) = 3*(3x^2+4)
3*f(-x) = 3*(3x^2)+3(4)
3*f(-x) = 9x^2+12